OpenStudy (anonymous):

What is the solution to the equation? 2(x – 7) = 10x + 18 @tHe_FiZiCx99

OpenStudy (the_fizicx99):



OpenStudy (anonymous):

thats 5

OpenStudy (the_fizicx99):

2x -14 = 10x +18 +14 +14 2x = 10x + 32 -10x -10 -8x = 32 /-8 /-8 x = -4

OpenStudy (anonymous):

what is 4/9 divided by 1/3

OpenStudy (the_fizicx99):

 Lol you try =.=